Prior to liquidating their partnership, Pepper and Haines had capital accounts of $18,000 and $65,000, respectively. The partnership assets were sold for $33,000. The partnership had no liabilities. Pepper and Haines share income and losses equally.

Required:

a. Determine the amount of Pepper's deficiency.
$

b. Determine the amount distributed to Haines, assuming Pepper is unable to satisfy the deficiency.
$
Answer #1

Loss on liquidation = Cash received - carrying value of assets

= 33,000 - (18,000+65,000)

= 50,000

Pepper's deficiency = Capital balance - share in loss on liquidation

= 18,000 - (50,000*1/2)

= 18,000 - 25,000

= 7,000

Amount distributed to Haines assuming pepper is unable to satisfy the deficiency

= 65,000 - 7,000 - 25,000

= 33,000